# Sheshathanks: Shesha or Sheshanaga is a serpentine demigod in Hindu mythology with multiple heads and serves as the celestial bed of Lord Vishnu.  : Multi-head Microarchitectural Leakage Discovery in new-generation Intel Processors

### v1

@@ Abstract
Transient execution attacks have been one of the widely explored microarchitectural side channels since the discovery of Spectre and Meltdown. However, much of the research has been driven by manual discovery of new transient paths through well-known speculative events. Although a few attempts exist in literature on automating transient leakage discovery, such tools focus on finding variants of known transient attacks and explore a small subset of instruction set. Further, they take a random fuzzing approach that does not scale as the complexity of search space increases. In this work, we identify that the search space of bad speculation is disjointedly fragmented into equivalence classes, and then use this observation to develop a framework named Shesha, inspired by Particle Swarm Optimization, which exhibits faster convergence rates than state-of-the-art fuzzing techniques for automatic discovery of transient execution attacks. We then use Shesha to explore the vast search space of extensions to the x86 Instruction Set Architecture (ISEs), thereby focusing on previously unexplored avenues of bad speculation. As such, we report five previously unreported transient execution paths in Instruction Set Extensions (ISEs) on new generation of Intel processors. We then perform extensive reverse engineering of each of the transient execution paths and provide root-cause analysis. Using the discovered transient execution paths, we develop attack building blocks to exhibit exploitable transient windows. Finally, we demonstrate data leakage from Fused Multiply-Add instructions through SIMD buffer and extract victim data from various cryptographic implementations.
@@ 1 Introduction
As per Intel, bad speculation happens due to three broader causes - branch misprediction, microcode assists and machine clear [23]. While a number of works have explored branch misprediction-based transient attacks in literature [10, 12, 30, 31, 35], very little research has been made to explore machine clears and microcode assists. Recently, authors in [41] explored the occurrences of machine clears in modern processors and demonstrated multiple transient execution paths based on different machine clear events. The findings of [41] open up a new frontier for transient execution discoveries that motivate further exploration. While most of the known works on transient leakage discoveries have been the results of manual effort, recent literature has seen quite some attempts to automate such leakage exploration [37, 21, 53, 22, 56]. However, these tools were developed to target certain classes of speculations and cannot be directly extended to explore the vast space of bad speculation (refer Sec. 9 for a detailed discussion and comparison with our tool). Additionally, due to the large number of x86 instructions, the prior works on automation only consider a small subset of Instruction Set Extensions (ISE), thereby missing out on a large pool of specialized instructions that are performance-bound and share various hardware resources. In fact, some of these ISEs, especially those concerned with SIMD, streaming SIMD, and Vector extensions, have their own separate data path (i.e. execution units), registers, temporal buffers, etc, making them an ideal spot for enabling speculation, and possibly bad speculation. Evidently, a systematic exploration of bad speculation in the context of ISEs is necessary to understand transient execution related vulnerabilities of ISEs. Concretely, in this work, we deal with the following question:
@@
The rest of the paper is organized as follows: Sec. 2 provides the necessary background information. Sec. 3, 4 and 5 are dedicated to the Shesha framework and its discoveries. In Sec. 6, we perform a root-cause analysis of the discovered transient execution paths and then discuss how they can aid as building blocks to micro-architectural attacks in Sec. 7. In Sec. 8, we show practical exploitation of leakage from FMA instructions. We then provide a discussion on related works in Sec. 9 and mitigation strategies in Sec. 10. Finally, we conclude in Sec. 11.
Responsible Disclosure: We have disclosed the results of our paper and the leakage due to FMA execution unit to Intel.
Artifact Availability: We will open-source the tool, datasets and proof-of-concept attack codes, post acceptance.
@@ 3 Shesha: Automatic Analysis of Bad Speculation in Instruction Set Extensions
In this section, we put forward a Particle Swarm Optimizer (PSO) inspired framework to perform an optimized exploration of search space related to bad speculation in ISEs. Towards this, we first elaborate on how the search space of bad speculation is disjointedly fragmented, allowing us to make modifications to generic PSO semantics (cf. Sec. 2.2) that provide faster convergence rates. Concretely, following the semantics of Sec. 2.2, instead of relying upon the entire swarm for swarm behaviour, the algorithm need only rely upon fragmented portions of the swarm, thereby allowing more granularity and enabling faster convergence. The disjointedly fragmented nature of the search space allows splitting Shesha into two phases: 1 Cognitive phase (where there is no social acceleration), and 2 Mixed phase (partial social/cognitive acceleration). By end of phase 1, each particle of the swarm is present in one of the disjointedly fragmented sub-space, and has knowledge of other particles in the same subspace. Thereby, in phase 2, the social acceleration component can be focused only on the fragmented sub-space. We detail this modification to the PSO semantics further in this section, along with the threat model.
@@ 3.2 Fitness function
As discussed in Sec. 2.2, the objective of a PSO is to optimize a well-defined fitness function $f:\mathcal{R}^{n}\rightarrow\mathcal{R}$. Since our search space is instruction sequences enabling bad speculation over ISEs, our fitness function must be a combination of indicators that capture the existence or absence of bad speculation. For a code sequence $\mathcal{C}$, in our context, $f$ can be defined as $f:\mathcal{C}\rightarrow\{0,1\}$, or $f$ provides a binary decision whether bad speculation was observed for a given code sequence $\mathcal{C}$. We defer discussion on the generation of $\mathcal{C}$ to Sec. 3.4.1 and focus on the binary decision of bad speculation here.
@@
Given this context, to implement the fitness function $f:\mathcal{C}\rightarrow\{0,1\}$, we require an interface to capture when any of 1 branch mis-prediction, 2 microcode assist, or 3 machine clear happens. All modern Intel systems expose a Performance Monitoring Unit [25] that logs different types of events occurring on the physical cores for analysis. Such performance counters are useful in understanding occurrences of bad speculation, given any code sequence $\mathcal{C}$. Tab. 1 summarizes the exact performance counters we used for Shesha. Therefore, given a code sequence $\mathcal{C}$, the objective of the PSO will be to maximize occurrences of bad speculation as reported by the respective performance counters.
@@ 3.3 Disjointedly Fragmented Sub-spaces
We now discuss the nature of bad speculation events that we term as disjointedly fragmented sub-spaces which allows us to achieve faster convergence for Shesha and efficiently scale as the number of ISE instruction pool increases. As evident from Tab. 1, each of the performance counters pertains to bad speculation scenarios, which are disjoint from each other from a causal perspective. For instance, bad speculation due to memory disambiguation issues has completely different cause (incorrect speculation while resolving load/store addresses) than bad speculation due to self-modifying code execution (incorrect speculation in writing to instruction cache). Additionally, the disjoint classes of bad speculation also have effects on disjoint architectural and micro-architectural assets. For example, memory disambiguation issues will mainly cause speculation in data forwarding in load/store buffers. On the other hand, self-modifying code execution issues will affect the instruction cache. Additionally, AVX-SSE transition issues shall affect the SIMD registers and temporal buffers. Therefore, we make the following observation:
Equivalence Classes: Different scenarios of bad speculation in Tab. 1 are disjoint wrt. cause and effect (on architectural/micro-architectural assets), thereby splitting the search space of bad speculation into disjointedly, fragmented equivalence classes, where the cause/effect of one equivalence class are unaffected by that of other classes.
Such a definition of equivalence classes (over set of instructions) ensures that the search space for bad speculation is split into disjointedly fragmented sub-spaces, such that the portion of the PSO’s swarm working on one equivalence class need not rely upon the entirety of the PSO swarm for exploitation (cf. Sec. 2.2). For example, without loss of generality, consider a portion of the PSO swarm working with machine clears related to self-modifying code. Then, the velocity vector computations related to swarm behaviour (i.e. involving $\mathbf{p}_{g}(i)$, cf. Sec. 2.2) need not consider the portion of the swarm working with bad speculation due to branch mispredictions, simply because code sequences triggering branch misprediction (for instance, heavy use of loops to train the branch predictor) have no direct bearing with code sequences triggering self-modifying code (for instance, use of store operations in address ranges belonging to .text section of the process binary, thereby requiring write operations on the instruction cache). Based on this description of equivalence classes, we now detail the construction of Shesha in the next subsection.
@@ 3.4 Swarm Optimizer description
With the idea of equivalence classes established, we now detail Shesha - an automatic analysis framework for exploring the disjointedly fragmented equivalence classes of bad speculation in ISEs. Shesha is divided mainly into three phases: 1 Swarm initialization, 2 Cognitive phase, and 3 Mixed phase. In phase 1, a population pool of N particles is created where each particle represents a randomly sampled instruction sequence from the instruction pools of different ISEs. Then, in phase 2, the particles are evolved to settle into sub-swarms (i.e. one swarm for one equivalence class). In this phase, since the sub-swarms are still being created, there is no exploitative behaviour (i.e. all particles evolve independently of each other). This design decision is intentional, following our discussion on equivalence classes in Sec. 3.3. Until the sub-swarms are ready, Shesha does not have reliable global best solutions to use for updating particle velocity (cf. Sec. 2.2 for PSO semantics). Finally, upon end of phase 2, the sub-swarms would be ready for each of the equivalence classes, which is when phase 3 kicks in and uses both local best and global best solutions to update particle velocity. We now explain these phases in detail.
@@ 3.4.1 Swarm Initialization
The first phase initializes a population of $N$ particles, where each represents an instruction sequence, sampled from the ISE instruction set. Following the semantics from Sec. 2.2, each particle is represented by a position vector $\mathbf{x}$ in a $n$-dimensional real-valued vector space, such that the fitness function $f:\mathcal{R}^{n}\rightarrow\mathcal{R}$ can be evaluated upon the same. We describe now how we map a code sequence to a real-valued $n$-dimensional vector.
@@ 3.4.2 Cognitive phase
In this phase, we allow for the compartmentalization of the swarm population into the disjointedly fragmented equivalence classes. Since the initialization phase has already computed the set of position vectors $\mathbb{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{N}\}$, we focus on generation of velocity vectors that allow generation of sub-swarms (i.e. one sub-swarm for each equivalence class). In order to do this, we must prevent any exploitation/swarm behaviour (cf. Sec. 2.2) from kicking in the cognitive phase. For this case, we choose $\alpha=1$, cognitive acceleration coefficient $\beta$ as $0.4$, and social acceleration coefficient $\gamma$ as 00 (to cancel out any swarm behaviour). This forces using the local best solutions $\mathbf{p}$ with probability $\beta$. In this phase, we rely upon two kinds of mutations to the velocity vector:
@@
    Modification to operands: Given a particle position vector $\mathbf{x}_{j}:1\leq j\leq N$, we sample a random dimension $d:1\leq d\leq n$ and mutate the operands with probability $\beta$. Following the semantics of 3.4.1, assume $x_{j}^{d}=$ (opcode << i) | operand_identifier_bitstring), where | represents Logical OR. Thereby, the new position vector is evaluated as $\mathbf{x}_{j}^{\prime}=\{x_{j}^{1},x_{j}^{2},...,x_{j}^{d-1},(x_{j}^{d}\,\wedge\,$ (1 << i)) << i | new_operand_identifier_bitstring, $x_{j}^{d+1},...,x_{j}^{n}\}$. Here, the newer operands that are sampled are converted to new_operand_identifier_bitstring using the same semantics as in Sec. 3.4.1. To follow upon the same example of VPXOR xmm1, xmm2, xmm3 having representation (vpxor_opcode << 12) | (1 << 8) | (2 << 4) | 3) in the position vector, assume the operand mutation converted this instruction to VPXOR xmm5, xmm13, xmm15. Then the modified representation of the aforementioned instruction in the position vector can be given by (vpxor_opcode << 12) | (5 << 8) | (13 << 4) | 15).
@@
Finally, the mutated position vector is evaluated upon the fitness function as defined in Sec. 3.2 to understand if the mutated particle discovers some newer avenues of bad speculation. As iterations of the Cognitive phase increase, the entire population of particles converges upon one of the several disjointedly fragmented equivalence classes. Thereby, at the end of the Cognitive phase, the sub-swarms belonging to each equivalence class $\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{9}$ are output to phase 2 (where each $S_{i}$ belongs to one equivalence class as in Tab. 1). Concretely, each particle in the initialized swarm belongs to exactly one of these equivalence classes.
@@ 3.4.3 Mixed phase
    Modifications to instruction: For each particle position vector $\mathbf{x}_{j}$ belonging to some sub-swarm $\mathcal{S}_{k}$, sample two random dimensions $d_{1},d_{2}:1\leq d_{1},d_{2}\leq n$ and mutate $x_{j}^{d_{1}}$ as in Cognitive phase with probability $\beta=0.1$. For dimension $d_{2}$, with probability $\gamma=0.4$, replace the $d_{2}$-th dimension of $\mathbf{x}_{j}$ with the $d_{2}$-th dimension of the leader of sub-swarm $S_{k}$. Here, leader of the sub-swarm $\mathcal{S}_{k}$ is defined as the position vector in sub-swarm $S_{k}$ with maximum fitness as described in Sec. 3.2 (i.e. the instruction sequence with maximum bad speculation in the equivalence class $\mathcal{S}_{k}$).
@@
At the end of this phase, Shesha outputs the population position vectors with maximized fitness (i.e. maximum occurrence of bad speculation) in the respective equivalence classes. We note that dimensionality reduction, in particular, ensures that the final output of Shesha is able to drop instructions from the generated instruction sequence that may not contribute to bad speculation. In other words, because of dimensionality reduction, the final output of Shesha is a minimal reproduced instruction sequence responsible for the relevant bad speculation scenarios detailed in Tab. 1.

### v2

@@ Abstract
Transient execution attacks have been one of the widely explored microarchitectural side channels since the discovery of Spectre and Meltdown. However, much of the research has been driven by manual discovery of new transient paths through well-known speculative events. Although a few attempts exist in literature on automating transient leakage discovery, such tools focus on finding variants of known transient attacks and explore a small subset of instruction set. Further, they take a random fuzzing approach that does not scale as the complexity of search space increases. In this work, we identify that the search space of bad speculation is disjointedly fragmented into equivalence classes, and then use this observation to develop a framework named Shesha, inspired by Particle Swarm Optimization, which exhibits faster convergence rates than state-of-the-art fuzzing techniques for automatic discovery of transient execution attacks. We then use Shesha to explore the vast search space of extensions to the x86 Instruction Set Architecture (ISAs), thereby focusing on previously unexplored avenues of bad speculation. As such, we report five previously unreported transient execution paths in Instruction Set Extensions (ISEs) on new generation of Intel processors. We then perform extensive reverse engineering of each of the transient execution paths and provide root-cause analysis. Using the discovered transient execution paths, we develop attack building blocks to exhibit exploitable transient windows. Finally, we demonstrate data leakage from Fused Multiply-Add instructions through SIMD buffer and extract victim data from various cryptographic implementations.
@@ 1 Introduction
As per Intel, bad speculation happens due to three broader causes - branch misprediction, microcode assists and machine clear [24]. While a number of works have explored branch misprediction-based transient attacks in literature [10, 12, 31, 32, 36], very little research has been made to explore machine clears and microcode assists. Recently, authors in [42] explored the occurrences of machine clears in modern processors and demonstrated multiple transient execution paths based on different machine clear events. The findings of [42] open up a new frontier for transient execution discoveries that motivate further exploration. While most of the known works on transient leakage discoveries have been the results of manual effort, recent literature has seen quite some attempts to automate such leakage exploration [38, 21, 54, 22, 57]. However, these tools were developed to target certain classes of speculations and cannot be directly extended to explore the vast space of bad speculation (refer Section 9 for a detailed discussion and comparison with our tool). Additionally, due to the large number of x86 instructions, the prior works on automation only consider a small subset of Instruction Set Extensions (ISE), thereby missing out on a large pool of specialized instructions that are performance-bound and share various hardware resources. In fact, some of these ISEs, especially those concerned with SIMD, streaming SIMD, and Vector extensions, have their own separate data path (i.e. execution units), registers, temporal buffers, etc, making them an ideal spot for enabling speculation, and possibly bad speculation. Evidently, a systematic exploration of bad speculation in the context of ISEs is necessary to understand transient execution related vulnerabilities of ISEs. Concretely, in this work, we deal with the following question:
@@
The rest of the paper is organized as follows: Section 2 provides the necessary background information. Section 3, 4 and 5 are dedicated to the Shesha framework and its discoveries. In Section 6, we perform a root-cause analysis of the discovered transient execution paths and then discuss how they can aid as building blocks to micro-architectural attacks in Section 7. In Section 8, we show practical exploitation of leakage from FMA instructions. We then provide a discussion on related works in Section 9 and mitigation strategies in Section 10. Finally, we conclude in Section 11.
Responsible Disclosure: We have disclosed the results of this work, including the leakage due to FMA execution unit to Intel. Intel acknowledged the leakage from the FMA execution unit, and responded that mitigation from [26] could cover this issue as well; therefore, no embargo period is required.
Artifact Availability: The tool and proof-of-concept attack codes can be found at https://github.com/SEAL-IIT-KGP/shesha.
@@ 3 Shesha: Automatic Analysis of Bad Speculation in Instruction Set Extensions
In this section, we put forward a Particle Swarm Optimizer (PSO) inspired framework to perform an optimized exploration of search space related to bad speculation in ISEs. Towards this, we first elaborate on how the search space of bad speculation is disjointedly fragmented, allowing us to make modifications to generic PSO semantics (cf. Section 2.2) that provide faster convergence rates. Concretely, following the semantics of Section 2.2, instead of relying upon the entire swarm for swarm behaviour, the algorithm need only rely upon fragmented portions of the swarm, thereby allowing more granularity and enabling faster convergence. The disjointedly fragmented nature of the search space allows splitting Shesha into two phases: 1 Cognitive phase (where there is no social acceleration), and 2 Mixed phase (partial social/cognitive acceleration). By end of phase 1, each particle of the swarm is present in one of the disjointedly fragmented sub-space, and has knowledge of other particles in the same subspace. Thereby, in phase 2, the social acceleration component can be focused only on the fragmented sub-space. We detail this modification to the PSO semantics further in this section, along with the threat model.
@@ 3.2 Fitness function
As discussed in Section 2.2, the objective of a PSO is to optimize a well-defined fitness function $f:\mathcal{R}^{n}\rightarrow\mathcal{R}$. Since our search space is instruction sequences enabling bad speculation over ISEs, our fitness function must be a combination of indicators that capture the existence or absence of bad speculation. For a code sequence $\mathcal{C}$, in our context, $f$ can be defined as $f:\mathcal{C}\rightarrow\{0,1\}$, or $f$ provides a binary decision whether bad speculation was observed for a given code sequence $\mathcal{C}$. We defer discussion on the generation of $\mathcal{C}$ to Section 3.4.1 and focus on the binary decision of bad speculation here.
@@
Given this context, to implement the fitness function $f:\mathcal{C}\rightarrow\{0,1\}$, we require an interface to capture when any of 1 branch mis-prediction, 2 microcode assist, or 3 machine clear happens. All modern Intel systems expose a Performance Monitoring Unit [25] that logs different types of events occurring on the physical cores for analysis. Such performance counters are useful in understanding occurrences of bad speculation, given any code sequence $\mathcal{C}$. Table 1 summarizes the exact performance counters we used for Shesha. Therefore, given a code sequence $\mathcal{C}$, the objective of the PSO will be to maximize occurrences of bad speculation as reported by the respective performance counters.
@@ 3.3 Disjointedly Fragmented Sub-spaces
We now discuss the nature of bad speculation events that we term as disjointedly fragmented sub-spaces which allows us to achieve faster convergence for Shesha and efficiently scale as the number of ISE instruction pool increases. As evident from Table 1, each of the performance counters pertains to bad speculation scenarios, which are disjoint from each other from a causal perspective. For instance, bad speculation due to memory disambiguation issues has completely different cause (incorrect speculation while resolving load/store addresses) than bad speculation due to self-modifying code execution (incorrect speculation in writing to instruction cache). Additionally, the disjoint classes of bad speculation also have effects on disjoint architectural and micro-architectural assets. For example, memory disambiguation issues will mainly cause speculation in data forwarding in load/store buffers. On the other hand, self-modifying code execution issues will affect the instruction cache. Additionally, AVX-SSE transition issues shall affect the SIMD registers and temporal buffers. Therefore, we make the following observation:
Equivalence Classes: Different scenarios of bad speculation in Table 1 are disjoint wrt. cause and effect (on architectural/micro-architectural assets), thereby splitting the search space of bad speculation into disjointedly, fragmented equivalence classes, where the cause/effect of one equivalence class are unaffected by that of other classes.
Such a definition of equivalence classes (over set of instructions) ensures that the search space for bad speculation is split into disjointedly fragmented sub-spaces, such that the portion of the PSO’s swarm working on one equivalence class need not rely upon the entirety of the PSO swarm for exploitation (cf. Section 2.2). For example, without loss of generality, consider a portion of the PSO swarm working with machine clears related to self-modifying code. Then, the velocity vector computations related to swarm behaviour (i.e. involving $\mathbf{p}_{g}(i)$, cf. Section 2.2) need not consider the portion of the swarm working with bad speculation due to branch mispredictions, simply because code sequences triggering branch misprediction (for instance, heavy use of loops to train the branch predictor) have no direct bearing with code sequences triggering self-modifying code (for instance, use of store operations in address ranges belonging to .text section of the process binary, thereby requiring write operations on the instruction cache). Based on this description of equivalence classes, we now detail the construction of Shesha in the next subsection.
@@ 3.4 Swarm Optimizer description
With the idea of equivalence classes established, we now detail Shesha - an automatic analysis framework for exploring the disjointedly fragmented equivalence classes of bad speculation in ISEs. Shesha is divided mainly into three phases: 1 Swarm initialization, 2 Cognitive phase, and 3 Mixed phase. In phase 1, a population pool of N particles is created where each particle represents a randomly sampled instruction sequence from the instruction pools of different ISEs. Then, in phase 2, the particles are evolved to settle into sub-swarms (i.e. one swarm for one equivalence class). In this phase, since the sub-swarms are still being created, there is no exploitative behaviour (i.e. all particles evolve independently of each other). This design decision is intentional, following our discussion on equivalence classes in Section 3.3. Until the sub-swarms are ready, Shesha does not have reliable global best solutions to use for updating particle velocity (cf. Section 2.2 for PSO semantics). Finally, upon end of phase 2, the sub-swarms would be ready for each of the equivalence classes, which is when phase 3 kicks in and uses both local best and global best solutions to update particle velocity. We now explain these phases in detail.
@@ 3.4.1 Swarm Initialization
The first phase initializes a population of $N$ particles, where each represents an instruction sequence, sampled from the ISE instruction set. Following the semantics from Section 2.2, each particle is represented by a position vector $\mathbf{x}$ in a $n$-dimensional real-valued vector space, such that the fitness function $f:\mathcal{R}^{n}\rightarrow\mathcal{R}$ can be evaluated upon the same. We describe now how we map a code sequence to a real-valued $n$-dimensional vector.
@@ 3.4.2 Cognitive phase
In this phase, we allow for the compartmentalization of the swarm population into the disjointedly fragmented equivalence classes. Since the initialization phase has already computed the set of position vectors $\mathbb{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{N}\}$, we focus on generation of velocity vectors that allow generation of sub-swarms (i.e. one sub-swarm for each equivalence class). In order to do this, we must prevent any exploitation/swarm behaviour (cf. Section 2.2) from kicking in the cognitive phase. For this case, we choose $\alpha=1$, cognitive acceleration coefficient $\beta$ as $0.4$, and social acceleration coefficient $\gamma$ as 00 (to cancel out any swarm behaviour). This forces using the local best solutions $\mathbf{p}$ with probability $\beta$. In this phase, we rely upon two kinds of mutations to the velocity vector:
@@
    Modification to operands: Given a particle position vector $\mathbf{x}_{j}:1\leq j\leq N$, we sample a random dimension $d:1\leq d\leq n$ and mutate the operands with probability $\beta$. Following the semantics of 3.4.1, assume $x_{j}^{d}=$ (opcode << i) | operand_identifier_bitstring), where | represents Logical OR. Thereby, the new position vector is evaluated as $\mathbf{x}_{j}^{\prime}=\{x_{j}^{1},x_{j}^{2},...,x_{j}^{d-1},(x_{j}^{d}\,\wedge\,$ (1 << i)) << i | new_operand_identifier_bitstring, $x_{j}^{d+1},...,x_{j}^{n}\}$. Here, the newer operands that are sampled are converted to new_operand_identifier_bitstring using the same semantics as in Section 3.4.1. To follow upon the same example of VPXOR xmm1, xmm2, xmm3 having representation (vpxor_opcode << 12) | (1 << 8) | (2 << 4) | 3) in the position vector, assume the operand mutation converted this instruction to VPXOR xmm5, xmm13, xmm15. Then the modified representation of the aforementioned instruction in the position vector can be given by (vpxor_opcode << 12) | (5 << 8) | (13 << 4) | 15).
@@
Finally, the mutated position vector is evaluated upon the fitness function as defined in Section 3.2 to understand if the mutated particle discovers some newer avenues of bad speculation. As iterations of the Cognitive phase increase, the entire population of particles converges upon one of the several disjointedly fragmented equivalence classes. Thereby, at the end of the Cognitive phase, the sub-swarms belonging to each equivalence class $\mathcal{S}_{1},\mathcal{S}_{2},...,\mathcal{S}_{9}$ are output to phase 2 (where each $S_{i}$ belongs to one equivalence class as in Table 1). Concretely, each particle in the initialized swarm belongs to exactly one of these equivalence classes.
@@ 3.4.3 Mixed phase
    Modifications to instruction: For each particle position vector $\mathbf{x}_{j}$ belonging to some sub-swarm $\mathcal{S}_{k}$, sample two random dimensions $d_{1},d_{2}:1\leq d_{1},d_{2}\leq n$ and mutate $x_{j}^{d_{1}}$ as in Cognitive phase with probability $\beta=0.1$. For dimension $d_{2}$, with probability $\gamma=0.4$, replace the $d_{2}$-th dimension of $\mathbf{x}_{j}$ with the $d_{2}$-th dimension of the leader of sub-swarm $S_{k}$. Here, leader of the sub-swarm $\mathcal{S}_{k}$ is defined as the position vector in sub-swarm $S_{k}$ with maximum fitness as described in Section 3.2 (i.e. the instruction sequence with maximum bad speculation in the equivalence class $\mathcal{S}_{k}$).
@@
At the end of this phase, Shesha outputs the population position vectors with maximized fitness (i.e. maximum occurrence of bad speculation) in the respective equivalence classes. We note that dimensionality reduction, in particular, ensures that the final output of Shesha is able to drop instructions from the generated instruction sequence that may not contribute to bad speculation. In other words, because of dimensionality reduction, the final output of Shesha is a minimal reproduced instruction sequence responsible for the relevant bad speculation scenarios detailed in Table 1.